# Externally Valid Policy Evaluation Combining Trial and Observational Data

### v2

@@ 4 Method
$\begin{split}\underline{W}^{\Gamma}&=\frac{1}{\Gamma}\cdot\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},\\ \overline{W}^{\Gamma}&=\Gamma\cdot\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},\end{split}$ (10)
@@
$c\cdot\frac{p(S=0|X)}{p(S=1|X)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},$ (11)
using Bayes’ rule. The bound equation 9 follows by applying equation 4. We proceed to show that the factors equation 10 are sufficient to construct an externally valid limit $\ell^{\Gamma}_{\alpha}$ for odds divergences up to degree $\Gamma$.
@@
$\widehat{F}(\ell;\mathcal{D}^{\prime\prime},w)=$ (13)
$\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell)}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell)+\overline{W}^{\Gamma}_{i}\mathds{1}(L_{i}>\ell)\big{]}+w},$ (14)
@@
$\boxed{\ell_{\alpha}^{\Gamma}(\mathcal{D})=\min_{\beta:0<\beta<\alpha}\ell_{\alpha,\beta}^{\Gamma},}$ (16)
@@ 5 Numerical Experiments
We will use both synthetic and real-world data to illustrate the main concepts of policy evaluation with limit curves $(\alpha,\ell^{\Gamma}_{\alpha})$. As a benchmark, we estimate the quantile using the inverse probability of sampling weighting Colnet et al. (2020)
@@
$\widehat{F}_{\text{IPSW}}(\ell;\mathcal{D})=$ (18)
$\frac{1}{m}\sum_{i=1}^{m}\frac{m}{n}\cdot\frac{\widehat{p}(S\!=\!0|X_{i})}{\widehat{p}(S\!=\!1|X_{i})}\cdot\frac{p_{\pi}(A_{i}|X_{i})}{p(A_{i})}\cdot\mathds{1}(L_{i}\leq\ell).$ (19)
@@ 5.1 Illustrations using synthetic data
$X|S=\begin{bmatrix}X_{0,S}\\ X_{1,S}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{0,S}\\ \mu_{1,S}\end{bmatrix},\begin{bmatrix}\sigma^{2}_{0,S}&0\\ 0&\sigma^{2}_{1,S}\end{bmatrix}\right),$ (21)
@@
where the parameters are given in Table 1. The distributions for populations A, B, C and Trial are taken to be unknown.
Table 1: Means and variances of covariate distribution $p(X|S)$ in equation 21.
Population $\mu_{0,S}$ $\mu_{1,S}$ $\sigma_{0,S}^{2}$ $\sigma_{1,S}^{2}$
A $(S=0)$ 0.5 0.5 1.0 1.0
B $(S=0)$ 0.0 0.0 1.5 1.5
C $(S=0)$ 0.5 0.5 1.5 1.5
Trial $(S=1)$ 0.0 0.0 1.0 1.0
The actions are binary $A\in\{0,1\}$ and corresponds to ‘do not treat’ versus ‘treat’. We begin by evaluating the ‘treat all’ policy, i.e.,
$p_{\pi_{1}}(A=1|X)=1.$ (22)
@@
The trial is an rct with equal probability of assignment, i.e., $p(A)\equiv 0.5$. The unknown conditional loss distribution is given by
$L|(A,X)$ $\sim\mathcal{N}(A\cdot X_{0}^{2}+X_{1}+2\cdot(1-A),1).$ (23)
The sampling probability $p(S|X)$ is treated as unknown. To illustrate the generality of the proposed methodology, we consider two different fitted models $\widehat{p}(S|X)$: a logistic model, which is conventionally used in the causal inference literature (Westreich et al., 2017), and the more flexible tree-based ensemble model trained by xgboost (Chen & Guestrin, 2016). The complete set of hyperparameters used is provided in the supplementary material.
@@
For the observational data $\big{(}X_{i}\big{)}^{n}_{i=1}$, we drew $n=2000$ samples. For the trial data we drew $1000$ samples: $m=500$ samples $\big{(}(X_{i},A_{i},L_{i})\big{)}^{m}_{i=1}$ were used to compute the limit curves, and the remaining $500$ samples were used to train $\widehat{p}(S|X)$.
(a) Sampling odds for target population A.
(b) Sampling odds for for target population B.
@@
Figure 4: True sampling odds $p(S=0|X)/p(S=1|X)$ compared with nominal odds obtained from logistic and xgboost models $\widehat{p}(S|X)$. The dots are a random subsample of the trial samples.
@@
Figure 3(a) compares nominal sampling odds obtained from the fitted models with the unknown odds $p(S=0|X)/p(S=1|X)$ for target population A. In this case, the logistic model happens to be well-specified so that the learned odds approximate the true ones well. The more flexible xgboost model also provide visually similar odds, albeit less accurate. By contrast, Figure 3(b) repeats the same exercise for target population B. Here the logistic model is misspecified and severely miscalibrated while the xgboost model continues to provide visually similar odds. A reliability diagram is used in Figure 3 for a quantitative metric for the divergence of the odds for the same scenarios. The reliability diagram uses 5 bins and shows the observed odds against the average predicted nominal odds $\widehat{p}(S=0|X)/\widehat{p}(S=1|X)$ for respective model. For target population A, both models are close to the diagonal and a degree of misspecification $\Gamma$ close to 1 is reasonable. For population B and the (misspecified) logistic model we see that a larger lower bound, $\Gamma\approx 1.5$, is appropriate, while for xgboost model $\Gamma\approx 1$ is still reasonable.
We then evaluate the out-of-sample loss of the ‘treat all’ policy $\pi_{1}$ using the limit curve for the benchmark and the limit curves for the proposed method. Figure 4(a) shows the evaluation with respect to population A, where all the limit curves are similar. The curves for the proposed method also illustrate increasing the credibility of the models results in less informative inferences. However, their informativeness stays above 90% for odds miscalibration degrees $\Gamma\in[1,1.5]$. Since the data is simulated, we evaluate the miscoverage gap of the curves in Figure 4(b). The gap is estimated using 1000 independent runs and for each run drawing $1000$ independent new samples $(X_{n+1},A_{n+1},L_{n+1})$. We see that the benchmark and the well-specified logistic model yields a valid and tight limit curve for $\Gamma=1$, whereas the xgboost model has only a slight negative miscoverage gap. As the degree of miscalibration $\Gamma$ increases to 1.5, the limit curves exhibit positive miscoverage gaps, where xgboost results in slightly less conservative inferences than the logistic model does.
@@
Figure 5: Evaluating ‘treat all’ policy $\pi_{1}$ for target population A with degrees of miscalibration $\Gamma\in[1,1.5]$.
We contrast the above results with the evaluation of $\pi_{1}$ when applied to population B as shown in Figure 5(a). The benchmark and the logistic model result in a similar limit curve, while xgboost infer higher losses of $\pi_{1}$. Figure 5(b) shows that the benchmark and the (misspecified) logistic model when assuming $\Gamma=1$ now yields negative miscoverage gaps, whereas xgboost yields a valid and tight limit curve. Again, when the degree of miscalibration increases to 1.5 both proposed models yield positive gaps. Our contrast between populations A and B shows that the use of more adaptable and, thus potentially better calibrated, models $\widehat{p}(S|X)$ can be advantageous over the conventional logistic or generalized linear models.
Figure 6: Evaluating ‘treat all’ policy $\pi_{1}$ for target population B with degrees of miscalibration $\Gamma\in[1,1.5]$.
We now turn to comparing the ‘treat all’ policy with a ‘treat none’ policy, i.e., $p_{\pi_{0}}(A=0|X)=1$, for population A. Their expected losses are estimated using equation 7 as $V^{\pi_{0}}_{\text{IPSW}}=2.73$ and $V^{\pi_{1}}_{\text{IPSW}}=1.85$, respectively. This evaluation suggests that $\pi_{1}$ is preferable to $\pi_{0}$. However, the limit curves presented in Figure 7 provide a more detailed picture in terms of out-of-sample losses: the tail losses certified for the ‘treat all’ policy $\pi_{0}$ are lower than those certified for $\pi_{1}$. This illustrates the cautionary principle built into the policy evaluations.
@@
The evaluation of $\pi_{1}$ with respect to population C for the logistic model of $\widehat{p}(S|X)$ is shown in Figure 1. More results for population C and results for additional populations are available in the supplementary material.
Figure 7: Limit curves for $\pi_{0}$ and $\pi_{1}$ for target population A certified for $\Gamma\in[1,1.5]$.
@@ 5.2 Evaluating seafood consumption policies
To illustrate the application of policy evaluation with real data, we study the impact of seafood consumption on blood mercury levels with data from the 2013-2014 National Health and Nutrition Examination Survey (NHANES). Following Zhao et al. (2019), each individual’s data includes eight covariates, denoted $X$, encompassing gender, age, income, the presence or absence of income information, race, education, smoking history, and the number of cigarettes smoked in the last month. We excluded one individual due to missing education data and seven individuals with incomplete smoking data. We impute the median income for 175 individuals with no income information. After the preprocessing, our data set comprises 1107 individuals. The data is then split into observational data $\mathcal{D}_{0}$ and trial data $\mathcal{D}_{1}$ with a probability $p(S=1|X_{\text{age}}\leq 35)=0.8$ respective $p(S=1|X_{\text{age}}>35)=0.2$ resulting in $637$ samples in the observational data and $470$ samples in the trial data. The action $A$ describes individual fish or shellfish consumption, categorizing an individual as having either low ($\leq$1 serving in the past month) or high ($>$12 servings in the past month) consumption. The loss $L$ represents the total concentration of blood mercury (measured in $\mu$g/L).
To generate counterfactual actions and losses, we consider a balanced rct so that $p(A)\equiv 0.5$ and learn a model of $p(L|A,X)$ from the data using gradient boosting (Freund & Schapire, 1997; Friedman et al., 2000). Thus during training, the trial data consists of samples $(X_{i},A_{i},L_{i})$ whereas the observational data only contains $X$.
A reliability diagram, as shown in Figure 8, compares a logistic model and a xgboost-trained model. The xgboost model is closer to the diagonal, and is therefore chosen and a $\Gamma$ close to $1$ is a reasonable lower limit.
Figure 8: Reliability diagram of the observed odds against the average predicted nominal odds obtained from logistic and xgboost-trained models.
In Figure 9 we compare the limit curve for a policy $\pi_{0}$ corresponding to low $(A\equiv 0)$ seafood consumption with the limit curve for a policy $\pi_{1}$ corresponding to high $(A\equiv 1)$ seafood consumption. We use an xgboost-trained model $\widehat{p}(S|X)$. For reference, a mercury level of 8 $\mu$g/L is guidance limit for women of child-bearing age. We see that under a low consumption policy a lower mercury level can be certified for miscalibrated odds $\Gamma\in[1,2]$. In this case, we can infer a 90% probability that the out-of-sample mercury level falls below the reference value of 8 $\mu$g/L.
Figure 9: Inferred blood mercury levels [$\mu$g/L] in a target population under ‘high’ and ‘low’ seafood consumption ($\pi_{1}$ and $\pi_{0}$, respectively). Limit curves for degrees of odds miscalibration $\Gamma\in[1,2]$.
@@ 6 Discussion
We have proposed a method for establishing externally valid policy evaluation based on experimental results from trial studies. The method is nonparametric, making no assumptions on the distributional forms of the data. Using covariate-only data from a target population, it takes into account possible covariate shifts between target and trial populations, and certifies valid finite-sample inferences of the out-of-sample loss $L_{n+1}$, up to any specified degree of model miscalibration.
@@
We also exemplify how the reliability diagram technique can be used to assess the divergences of the model odds and facilitate a more systematic guidance on both model selection and on the specification of the odds miscalibration degree $\Gamma$ in a given problem.
7 Broader Impact
The method we propose is designed for safety-critical applications, such as clinical decision support, with the cautionary principle in mind. We believe that it offers a valuable tool for policy evaluation in such scenarios. Our approach focuses on limit curves, coupled with a statistical guarantee, for a more detailed understanding of the out-of-sample loss. This facilitates a fairer evaluation by bringing attention to sensitive covariates in the tails. However, it remains important to be aware of biases, and it might be necessary to address them separately to prevent their replication.
@@ Appendix A Proof
The proof technique presented here builds upon several results established in Vovk et al. (2005); Tibshirani et al. (2019); Jin et al. (2023); Ek et al. (2023).
Use equation 15 to construct the limit
@@
for any $0<\beta<\alpha$, where $\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$ is defined in equation 12. We want to lower bound the probability of $L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D})$. First, note that
@@
whenever $\overline{W}_{n+1}^{\Gamma}\leq\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$. The second line follows from using sample splitting, which ensures that $L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})$ and $\overline{W}_{n+1}^{\Gamma}\leq\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$ are independent events.
To lower bound $\mathbb{P}\{L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})\}$, we will make use of the following inquality,
$\mathbb{E}\left[\widehat{F}(\ell_{\alpha}^{\Gamma};\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})\right]=\mathbb{E}\left[\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}}\right]\geq\frac{1-\alpha}{1-\beta},$ (31)
@@
$\big{(}(X_{m^{\prime}_{0}+1},A_{m^{\prime}_{0}+1},L_{m^{\prime}_{0}+1}),\dots,(X_{m},A_{m},L_{m}),(X_{n+1},A_{n+1},L_{n+1})\big{)}.$ (32)
From Tibshirani et al. (2019) we have that the out-of-sample loss $L_{n+1}$ has the (conditional) cdf
$\mathbb{P}\{L_{n+1}\leq\ell\;|\;\mathcal{S}_{+}\}=\sum_{i\in\mathcal{S}_{+}}p_{i}\mathds{1}(\ell_{i}\leq\ell)=\frac{\sum_{i\in\mathcal{S}_{+}}w_{i}\mathds{1}(L_{i}\leq\ell)}{\sum_{i\in\mathcal{S}_{+}}w_{i}}.$ (33)
Next, we build on the proof method used in Jin et al. (2023, thm. 2.2). Use the limit $\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})$ from equation 28 in equation 33 and apply the law of total expectation to perform marginalization over $\mathcal{S}_{+}$
@@
$\geq\mathbb{E}\left[\frac{\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{S}_{+}}W_{i}}\right]-\mathbb{E}\left[\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}}\right]$ (37)
$=\mathbb{E}\left[\frac{(*)}{\left[\sum_{i\in\mathcal{S}_{+}}W_{i}\right]\left[\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}\right]}\right],$ (38)

### v3

@@ 4.2 Inferring out-of-sample limit
$\underline{W}^{\Gamma}=\frac{1}{\Gamma}\cdot\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},\qquad\overline{W}^{\Gamma}=\Gamma\cdot\frac{\widehat{p}(S=0|X)}{\widehat{p}(S=1|X)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},$ (11)
@@
$c\cdot\frac{p(S=0|X,U)}{p(S=1|X,U)}\cdot\frac{p_{\pi}(A|X)}{p(A|X)},$ (12)
using Bayes’ rule. The bound equation 10 follows by applying equation 3. We proceed to show that the factors equation 11 are sufficient to construct an externally valid limit $\ell^{\Gamma}_{\alpha}$ for odds divergences up to degree $\Gamma$, similar to Ek et al. (2023).
@@
$\widehat{F}(\ell;\mathcal{D}^{\prime\prime},w)=\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell)}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell)+\overline{W}^{\Gamma}_{i}\mathds{1}(L_{i}>\ell)\big{]}+w},$ (14)
@@
$\ell_{\alpha}^{\Gamma}(\mathcal{D})=\min_{\beta:0<\beta<\alpha}\ell_{\alpha,\beta}^{\Gamma},$ (16)
@@
Figure 4: Reliability diagram of the observed odds against the average predicted nominal odds obtained from models $\widehat{p}(S|X)$.
@@ 5 Numerical Experiments
We will use both synthetic and real-world data to illustrate the main concepts of policy evaluation with limit curves $(\alpha,\ell^{\Gamma}_{\alpha})$. As a performance benchmark, we estimate the quantile – which yields the tightest limit – using the inverse probability of sampling weighting (Colnet et al., 2024)
@@
$\widehat{F}_{\text{IPSW}}(\ell;\mathcal{D})=\frac{1}{m}\sum_{i=1}^{m}\frac{m}{n}\cdot\frac{\widehat{p}(S\!=\!0|X_{i})}{\widehat{p}(S\!=\!1|X_{i})}\cdot\frac{p_{\pi}(A_{i}|X_{i})}{p(A_{i})}\cdot\mathds{1}(L_{i}\leq\ell).$ (18)
@@ 5.1 Illustrations using synthetic data
$X|S=\begin{bmatrix}X_{0,S}\\ X_{1,S}\end{bmatrix}\sim\mathcal{N}\left(\begin{bmatrix}\mu_{0,S}\\ \mu_{1,S}\end{bmatrix},\begin{bmatrix}\sigma^{2}_{0,S}&0\\ 0&\sigma^{2}_{1,S}\end{bmatrix}\right),\qquad U|S\sim\mathcal{N}(\mu_{U,S},\sigma^{2}_{U,S}),$ (20)
@@
where the parameters are given in Table 1. The distributions for populations A, B and Trial are taken to be unknown.
Table 1: Means and variances of covariate distribution $p(X,U|S)$ in equation 20.
Population $\mu_{0,S}$ $\mu_{1,S}$ $\mu_{U,S}$ $\sigma_{0,S}^{2}$ $\sigma_{1,S}^{2}$ $\sigma_{U,S}^{2}$
A $(S=0)$ 0.5 0.5 0.5 1.0 1.0 1.0
B $(S=0)$ 0.0 0.5 0.0 1.25 1.5 1.25
Trial $(S=1)$ 0.0 0.0 0.0 1.0 1.0 1.0
@@
The actions are binary $A\in\{0,1\}$ and corresponds to ‘do not treat’ versus ‘treat’. We evaluate the ‘treat all’ policy, i.e. $p_{\pi_{1}}(A=1|X)=1$. The trial is an rct with equal probability of assignment, i.e., $p(A)\equiv 0.5$. The unknown conditional loss distribution is given by
$L|(A,X,U)$ $\sim\mathcal{N}(A\cdot X_{0}^{2}+X_{1}+A\cdot U+(1-A),1).$ (21)
The sampling probability $p(S|X,U)$ is treated as unknown. The complete set of hyperparameters used is provided in the supplementary material.
@@
For the observational data $\big{(}X_{i}\big{)}^{n}_{i=1}$, we drew $n=2000$ samples. For the trial data we drew $1000$ samples: $m=500$ samples and $\big{(}(X_{i},A_{i},L_{i})\big{)}^{m}_{i=1}$ was used to compute the limit curves. The remaining $500$ samples were used to train $\widehat{p}(S|X)$.
(a) Selection odds for target population A.
(b) Selection odds for for target population B.
@@
Figure 5: Odds $p(S=0|X)/p(S=1|X)$ compared with nominal odds obtained from logistic and xgboost models $\widehat{p}(S|X)$. The dots are a random subsample of the trial samples.
@@
Figure 5(a) compares nominal selection odds obtained from the fitted models with the unknown odds $p(S=0|X)/p(S=1|X)$ for target population A. We consider two different fitted models $\widehat{p}(S|X)$: a logistic model, which is conventionally used in the causal inference literature (Westreich et al., 2017), and the more flexible tree-based ensemble model trained by xgboost (Chen and Guestrin, 2016). In this case, the logistic model happens to be well-specified so that the learned odds approximate the true ones well. The more flexible xgboost model also provide visually similar odds, albeit less accurate. By contrast, Figure 5(b) repeats the same exercise for target population B. Here the logistic model is misspecified and severely miscalibrated while the xgboost model continues to provide visually similar odds. A well-performing and flexible model is required for a meaningful benchmark of the upper value of $\Gamma$. Therefore, we will use the xgboost model.
In Figure 4, we use the reliability diagram technique to assess the performance of the xgboost model $\widehat{\text{odds}}(X)$ of the nominal odds for target population B. The model is close to the diagonal suggesting that it is sufficiently flexible to accurately model the odds. This is in line with the result in Figure 5(b). We then use observed covariates to calibrate appropriate upper bounds for $\Gamma$. Figure 6(a) shows the evaluation with respect to population B. Assuming that the unmeasured $U$ have no greater selection effect than $X_{0}$, a degree of miscalibration $\Gamma=2$ is a credible choice as it covers more than $95\%$ of the ratios. Since the data is simulated, we evaluate the miscoverage gap of the limit curves of the ‘treat all’ policy $\pi_{1}$ in Figure 6(b) for the benchmark and the limit curves for the proposed method. The gap is estimated using 1000 independent runs and for each run drawing $1000$ independent new samples $(X_{n+1},U_{n+1},A_{n+1},L_{n+1})$. We see that the benchmark and the limit curve for $\Gamma=1$ yields a negative miscoverage gap. As the degree of miscalibration $\Gamma$ increases to 2, the limit curves exhibit positive miscoverage gaps.
@@
Figure 6: Evaluating a ‘treat all’ policy $\pi_{1}$ for target population B. (a) Benchmarking the degree of miscalibration $\Gamma$ using omitted covariates. (b) Miscoverage gaps when degree of miscalibration $\Gamma\in[1,2]$.
@@
The evaluation of $\pi_{1}$ with respect to population A is shown in Figure 1. Results for the logistic model, another policy $\pi_{0}$, and for additional populations are available in the supplementary material.
@@ 5.2 Evaluating seafood consumption policies
To illustrate the application of policy evaluation with real data, we study the impact of seafood consumption on blood mercury levels with data from the 2013-2014 National Health and Nutrition Examination Survey (NHANES). Following Zhao et al. (2019), each individual’s data includes eight covariates, encompassing gender, age, income, the presence or absence of income information, race, education, smoking history, and the number of cigarettes smoked in the last month. All covariates, except for smoking history $U$, are treated as measured and denoted $X$. We excluded one individual due to missing education data and seven individuals with incomplete smoking data. We impute the median income for 175 individuals with no income information. The continuous covariates – age, income and number of cigarettes smoked in the last month – are standardized. After the preprocessing, our data set comprises 1107 individuals. The data is then split into observational data $\mathcal{D}_{0}$ and trial data $\mathcal{D}_{1}$ based on the covariates gender, age, income and smoking history (more details are available in the supplementary material) resulting in $646$ samples in the observational data and $461$ samples in the trial data. The action $A$ describes individual fish or shellfish consumption, categorizing an individual as having either low ($\leq$1 serving in the past month) or high ($>$12 servings in the past month) consumption. The loss $L$ represents the total concentration of blood mercury (measured in $\mu$g/L).
To generate counterfactual actions and losses, we consider a balanced rct so that $p(A)\equiv 0.5$ and learn a model of $p(L|A,X,U)$ from the data using gradient boosting (Freund and Schapire, 1997; Friedman et al., 2000). Thus during training, the trial data consists of samples $(X_{i},A_{i},L_{i})$ whereas the observational data only contains $X$.
In Figure 3(a) we use observed covariates to benchmark appropriate upper bounds for $\Gamma$. We exclude each of the seven covariates one by one, highlighting the three most dominant ones in the figure. If the unmeasured covariates have weaker influence than these, a $\Gamma$ value in the range of $1.5$ to $2$ is appropriate. In Figure 3(b) we compare the limit curve for a policy $\pi_{0}$ corresponding to low $(A\equiv 0)$ seafood consumption with the limit curve for a policy $\pi_{1}$ corresponding to high $(A\equiv 1)$ seafood consumption. We use an xgboost-trained model $\widehat{p}(S|X)$. For reference, a mercury level of 8 $\mu$g/L is guidance limit for women of child-bearing age. We see that under a low consumption policy a lower mercury level can be certified for miscalibrated odds $\Gamma\in[1,2]$. In this case, we can infer a 90% probability that the out-of-sample mercury level falls below the reference value of 8 $\mu$g/L.
@@ 6 Discussion
We have proposed a method for establishing externally valid policy evaluation based on experimental results from trial studies. The method is nonparametric, making no assumptions on the distributional forms of the data. Using additional covariate data from a target population, it takes into account possible covariate shifts between target and trial populations, and certifies valid finite-sample inferences of the out-of-sample loss $L_{n+1}$, up to any specified degree of model miscalibration.
@@
We also exemplify how the reliability diagram technique and the benchmark approach of omitting observed selection factors facilitate a more systematic guidance on the specification of the odds miscalibration degree $\Gamma$ in any given problem.
7 Broader Impact and Limitations
The method we propose is designed for safety-critical applications, such as clinical decision support, with the cautionary principle in mind. We believe that it offers a valuable tool for policy evaluation in such scenarios. Our approach focuses on limit curves, coupled with a statistical guarantee, for a more detailed understanding of the out-of-sample loss. This facilitates a fairer evaluation by bringing attention to sensitive covariates in the tails. However, it remains important to be aware of biases, and it might be necessary to address them separately to prevent their replication. It is also important to note that the method requires independent samples, a condition that may not be met during major virus outbreaks or similar situations.
Acknowledgments and Disclosure of Funding
This work was partially supported by the Wallenberg AI, Autonomous Systems and Software Program (WASP) funded by the Knut and Alice Wallenberg Foundation, and the Swedish Research Council under contract 2021-05022.
Supplementary Material
In this supplementary material, we provide the proof outlined in Section A and additional details on the numerical experiments discussed in Section B.
@@ Appendix A Proof
In this context, $\mathbb{P}$ represents the probability over samples drawn from both $p(X,U,A,L|S=1)$ and $p_{\pi}(X,U,A,L|S=0)$. The proof technique presented here builds upon several results established in Vovk et al. (2005); Tibshirani et al. (2019); Jin et al. (2023); Ek et al. (2023), and for completeness we present the proof in full.
Following Ek et al. (2023) we introduce $\beta$ such that $0<\beta<\alpha$. Use equation 15 to construct the limit
@@
where $\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$ is defined in equation 13. We want to lower bound the probability of $L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D})$. Note that
@@
whenever $\overline{W}_{n+1}^{\Gamma}\leq\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$. The second line results from applying sample splitting, which gurantess that $L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})$ and $\overline{W}_{n+1}^{\Gamma}\leq\overline{w}_{\beta}^{\Gamma}(\mathcal{D}^{\prime})$ are independent.
To lower bound $\mathbb{P}\{L_{n+1}\leq\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})\}$, we will make use of the following inequality,
$\mathbb{E}\left[\widehat{F}(\ell_{\alpha}^{\Gamma};\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})\right]=\mathbb{E}\left[\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\overline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}}\right]\geq\frac{1-\alpha}{1-\beta},$ (29)
@@
$\big{(}(X_{m^{\prime}+1},U_{m^{\prime}+1},A_{m^{\prime}+1},L_{m^{\prime}+1}),\dots,(X_{m},U_{m},A_{m},L_{m}),(X_{n+1},U_{n+1},A_{n+1},L_{n+1})\big{)}.$ (30)
From Tibshirani et al. (2019, thm. 2) we have that the out-of-sample loss $L_{n+1}$ has the (conditional) cdf
$\mathbb{P}\{L_{n+1}\leq\ell\;|\;\mathcal{S}_{+}\}=\sum_{i\in\mathcal{S}_{+}}p_{i}\mathds{1}(\ell_{i}\leq\ell)=\frac{\sum_{i\in\mathcal{S}_{+}}w_{i}\mathds{1}(L_{i}\leq\ell)}{\sum_{i\in\mathcal{S}_{+}}w_{i}},$ (31)
where $w_{i}$ quantifies the distribution shift for sample $i$ using the (unobservable) ratio equation 9. Next, we build on the proof method used in Jin et al. (2023, thm. 2.2). Use the limit $\ell_{\alpha}^{\Gamma}(\mathcal{D}^{\prime\prime},\overline{W}_{n+1}^{\Gamma})$ from equation 26 in equation 31 and apply the law of total expectation to perform marginalization over $\mathcal{S}_{+}$
@@
$\geq\mathbb{E}\left[\frac{\sum_{i\in\mathcal{S}_{+}}W_{i}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{S}_{+}}W_{i}}\right]-\mathbb{E}\left[\frac{\sum_{i\in\mathcal{D}^{\prime\prime}}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})}{\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\overline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}}\right]$ (35)
$=\mathbb{E}\left[\frac{(*)}{\left[\sum_{i\in\mathcal{S}_{+}}W_{i}\right]\left[\sum_{i\in\mathcal{D}^{\prime\prime}}\big{[}\underline{W}_{i}^{\Gamma}\mathds{1}(L_{i}\leq\ell_{\alpha}^{\Gamma})+\overline{W}_{i}^{\Gamma}\mathds{1}(L_{i}>\ell_{\alpha}^{\Gamma})\big{]}+\overline{W}_{n+1}^{\Gamma}\right]}\right],$ (36)